# SPEAR: A Unified SSL Framework for Learning Speech and Audio Representations

### v1

@@ 5.3 HEAR Evaluation
| Model | # Params | Pre-train Data | Env | Speech | Music | Average |
@@
| WavLM Base+ (Chen et al., 2022) | 95M | 94k | 57.28 | 68.14 | 61.31 | 62.69 |
| WavLM Large (Chen et al., 2022) | 317M | 94k | 72.86 | 72.69 | 65.77 | 69.65 |
@@
| Audio Models | ||||||
| BEATs (Chen et al., 2023) | 90M | 5k | 73.23 | 62.40 | 77.52 | 71.05 |
[…]
| Dasheng-Base (Dinkel et al., 2024) | 86M | 272k | 80.18 | 72.48 | 84.00 | 79.31 |
| Dasheng 0.6B (Dinkel et al., 2024) | 600M | 272k | 82.95 | 74.82 | 84.73 | 81.03 |
| Dasheng 1.2B (Dinkel et al., 2024) | 1.2B | 272k | 83.20 | 75.72 | 84.86 | 81.44 |
@@
[…]
| Ours, SPEAR Base | 94M | 97k | 80.66 | 73.73 | 79.29 | 77.75 |
@@
| Ours, SPEAR XLarge | 600M | 197k | 81.74 | 76.76 | 80.92 | 79.72 |
[…]
Setup
To assess the general audio capabilities of our models, experiments are conducted on the HEAR benchmark (Turian et al., 2022), which evaluates audio representations across 19 diverse tasks. The final-layer representations are used for evaluation unless otherwise specified. For clarity, the average scores for each of the three task categories are reported: Environment, Speech and Music, along with the overall average score in Table 6. More information regarding the tasks in HEAR and detailed results on individual tasks are provided in Appendix F.
Results
As shown in Table 6, all speech-domain models show very limited overall performance on the general-audio-focused HEAR benchmark, highlighting the need to incorporate general audio data during pre-training. Nonetheless, our SPEAR models, even the SPEAR Base, yield a higher overall score than the much bigger WavLM Large pre-trained with k-means tokens, indicating that the fine-grained discrete tokens manage to capture general-audio-related information despite being extracted from a speech SSL model. Our audio-domain models demonstrate strong performance on environment-related tasks. Notably, SPEAR Large achieves 83.58 on the environment category, surpassing the best performing audio SSL model Dasheng 1.2B (83.2), with only a quarter of the model parameters and far less pre-training data. However, its overall performance on speech and music tasks trails that of Dasheng 1.2B, a gap we attribute to the significantly smaller scale of SPEAR Large in terms of pre-training data and model size.
Finally, our dual-domain models consistently outperform their single-domain counterparts, highlighting the benefits of our unified pre-training framework. The SPEAR Large achieves an average score of 79.26, surpassing both SPEAR Large (72.96) and SPEAR Large (79.18), indicating that SPEAR successfully unifies the representation space of both domains through joint pre-training on both speech and audio data, making it a more versatile model.
@@ Results
Compared to USAD Large trained with multi-teacher knowledge distillation, our SPEAR Large pre-trained on a smaller corpora achieves a significant absolute improvement of 2.42 on the average score under the same evaluation setup of using intermediate representations. This again highlights the strength of SPEAR as a unified SSL framework for learning generic speech and audio representations jointly from speech and general audio data. Finally, compared to the much larger Dasheng 1.2B, SPEAR XLarge achieves stronger performance on speech-related tasks while trailing on environment and music tasks, which can be attributed to its smaller model size and imbalanced pre-training data composition (only 13k hours from the 197k hours are general audio data). However, this performance gap can be reversed by leveraging all intermediate layer representations.
@@ 6 Conclusions
In this work, we propose SPEAR, a unified SSL framework for both speech and general audio domains, learning unified and generic representations across both domains. By leveraging multi-codebook vector quantisation to generate fine-grained discrete speech and audio tokens, SPEAR performs fine-grained masked-token prediction as the pre-training task for representation learning. Based on this, an asymmetrical dual-domain pre-training pipeline is designed for balancing the performance on both domains. To the best of our knowledge, SPEAR is the first SSL framework to successfully learn unified speech and audio representations from a mixture of speech and audio. The downstream fine-tuning experiments, along with the evaluation of frozen representations on two major benchmarks for evaluating speech and general-audio representations (SUPERB and HEAR), demonstrate the effectiveness of SPEAR in learning unified and generic speech and audio representations. Our dual-domain model with 600M parameters excels in both domains, making it one of the most powerful and versatile open-source SSL models for auditory understanding.
7 Ethics Statement
@@
8 Reproducibility Statement
To ensure the reproducibility of our research and to support further advancements in the field, we will make our resources publicly available, including the inference code and the model checkpoints. All essential training details, including model configurations and hyperparameters, have been thoroughly documented in the main paper as well as Appendix C.2 and Appendix C.3. We hope that by providing these resources, more researchers can contribute to the development of this exciting research area. Details regarding access and implementation will be updated after the double-blind review. We invite the community to build upon our work to further advance the research field.
Appendix A LLM Usage
We used LLMs in paper presentation for the purpose of correcting grammatical errors and spelling.
Appendix B Multi-codebook Vector Quantisation
@@
B.1 Encode and Decode
A Multi-codebook Vector quantisation (MVQ) module consists of codebooks, each containing codebook vectors. Given a -dimensional representation , the MVQ quantiser encodes it to a sequence of integers (i.e, tokens) from a finite discrete value space .444For storage efficiency, we always use since the indices can be stored with uint8 format. However, it is theoretically possible to increase to a bigger number. The encoded integers are denoted as MVQ tokens, which can be used for reconstructing the original input through a operation:
@@ B.2 MVQ Training
Appendix C Model Specification and Training Settings
C.1 Model Specification
@@
C.2 Pre-training Configurations
The training configurations adopted in SPEAR are presented in Table 8. As shown in WavLM (Chen et al., 2022), applying data augmentations to the input audio can improve the pre-training performance. Therefore, we apply both in-batch utterance mixing and noise mixing during training. MUSAN (Snyder et al., 2015) is used as the noise dataset. The optimiser and scheduler settings follow Yao et al. (2024), where the ScaledAdam optimiser and Eden scheduler are used.
@@ C.2 Pre-training Configurations
| (see Equation 6) | - | - | - | - | 0.1 | 0.1 | 0.1 |
C.3 Fine-tuning Configurations
The fine-tuning configurations for the downstream ASR tasks and AT tasks presented in Section 5.1 are shown in Table 9. We used Pruned RNN-T (Kuang et al., 2022), a memory-efficient variant of RNN-T for optimisation. An asynchronous learning rate policy is adopted during fine-tuning by setting a smaller learning rate for the pre-trained encoder parameters. In ASR experiments, 3-fold speed perturbation is applied to the training data. In AS-2M, we follow prior work (Gong et al., 2021) to adopt a weighted sampler to cope with the imbalanced label distribution in the full AudioSet. Mixup (Zhang et al., 2018) with a probability of 0.5 is used in AT fine-tuning.
@@ C.3 Fine-tuning Configurations
Appendix D LibriSpeech Fine-tuning Experiments
@@
Appendix E SUPERB Evaluation
In this section, we provide more detail about the SUPERB benchmark (Yang et al., 2021; Tsai et al., 2022) as a supplement to Section 5.2 and present the complete results on SUPERB benchmark. A summary of the SUPERB tasks is shown in Table 11. Complete results of SPEAR models along with other existing speech SSL models are shown in Table 12 and Table 13.
| Task Name | Metric(s) | ||
| Speaker Identification (SID) | Accuracy | ||
| Automatic Speaker Verification (ASV) | Equal Error Rate (EER) | ||
| Speaker Diarization (SD) | Diarization error rate (DER) | ||
| Emotion Recognition (ER) | Accuracy | ||
| Phoneme Recognition (PR) | Phone Error Rate (PER) | ||
| Automatic Speech Recognition (ASR) | Word Error Rate (WER) | ||
| Out-of-domain ASR (ar/es/zh) | Word (character) Error Rate | ||
| Keyword Spotting (KS) | Accuracy | ||
| Query by Example Spoken Term Detection (QbE) | Maximum Term Weighted Value (MTWV) | ||
| Speech Translation (ST) | BLEU | ||
| Intent Classification (IC) | Accuracy | ||
| Slot Filling (SF) | F1, Character Error Rate (CER) | ||
| Speech Enhancement (SE) |
| Speech Separation (SS) | Scale-invariant Signal-to-distortion Ratio improvement (SI-SDRi) | ||
@@ Appendix E SUPERB Evaluation
| Model | # Params |
Understanding | ||||||||||
| PR | ASR | OOD-ASR | KS | QbE | ST | IC | SF | ||||||
| PER | WER | WER | Acc | MTWV | BLEU | Acc | F1 | CER | |||||
| FBANK | 0 | - | 82.01 | 23.18 | 63.58 | 8.63 | 0.0058 | 2.32 | 9.10 | 69.64 | 52.94 | ||
| Existing Speech SSL models | |||||||||||||
| WavLM Base+ (Chen et al., 2022) | 95M | 94k | 3.92 | 5.59 | 38.32 | 97.37 | 0.0988 | 24.25 | 99.00 | 90.58 | 21.20 | ||
| wav2vec 2.0 Large (Baevski et al., 2020) | 317M | 60k | 4.25 | 3.75 | 44.89 | 96.66 | 0.0480 | 12.48 | 95.28 | 87.11 | 27.31 | ||
| HuBERT Large (Hsu et al., 2021) | 317M | 60k | 3.53 | 3.62 | 44.08 | 95.29 | 0.0353 | 20.10 | 98.76 | 89.81 | 21.76 | ||
| WavLM Large (Chen et al., 2022) | 317M | 94k | 3.06 | 3.44 | 32.27 | 97.86 | 0.0886 | 26.57 | 99.31 | 92.21 | 18.36 | ||
| Ours, Speech SSL models | |||||||||||||
| SPEAR Base | 94M | 84k | 3.44 | 3.46 | 34.35 | 97.50 | 0.0772 | 24.37 | 99.17 | 90.96 | 19.22 | ||
| SPEAR Large | 327M | 84k | 2.56 | 3.27 | 31.70 | 97.89 | 0.0768 | 26.20 | 99.47 | 92.25 | 17.86 | ||
| Ours, Speech & Audio SSL models | |||||||||||||
| SPEAR Base | 94M | 97k | 3.89 | 3.76 | 35.48 | 97.58 | 0.0801 | 24.07 | 98.05 | 90.54 | 20.14 | ||
| SPEAR Large | 327M | 97k | 3.08 | 3.39 | 31.22 | 97.92 | 0.0712 | 25.64 | 99.40 | 92.07 | 18.04 | ||
| SPEAR XLarge | 600M | 197k | 2.94 | 3.19 | 30.69 | 98.12 | 0.0745 | 26.66 | 99.61 | 92.86 | 17.23 | ||
| Model | # Params |
Paralinguistics | Enhancement | Generation | |||||||||
| SID | ASV | SD | ER | SE | SS | VC | ||||||||
| Acc | EER | DER | Acc | PESQ | STOI | SI-SDRi | MCD | WER | ASV | |||||
| FBANK | 0 | - | 0 | 9.56 | 10.05 | 35.39 | 2.55 | 93.6 | 9.23 | 8.47 | 38.3 | 77.25 | ||
| Existing Speech SSL models | ||||||||||||||
| WavLM Base+ (Chen et al., 2022) | 95M | 94k | 89.42 | 4.07 | 3.50 | 68.65 | 2.63 | 94.3 | 10.85 | 7.40 | 8.1 | 99.00 | ||
| wav2vec 2.0 Large (Baevski et al., 2020) | 317M | 60k | 86.14 | 5.65 | 5.62 | 65.64 | 2.52 | 94.0 | 10.02 | 7.63 | 15.8 | 97.25 | ||
| HuBERT Large (Hsu et al., 2021) | 317M | 60k | 90.33 | 5.98 | 5.75 | 67.62 | 2.64 | 94.2 | 10.45 | 7.22 | 9.0 | 99.25 | ||
| WavLM Large (Chen et al., 2022) | 317M | 94k | 95.49 | 3.77 | 3.24 | 70.62 | 2.70 | 94.5 | 11.19 | 7.30 | 9.9 | 99.00 | ||
| Ours, Speech SSL models | ||||||||||||||
| SPEAR Base | 94M | 84k | 90.5 | 3.75 | 3.57 | 69.21 | 2.64 | 94.3 | 10.84 | 7.40 | 10.1 | 99.00 | ||
| SPEAR Large | 327M | 84k | 95.49 | 3.14 | 3.20 | 72.10 | 2.71 | 94.5 | 11.20 | 7.33 | 10.4 | 99.00 | ||
| Ours, Speech & Audio SSL models | ||||||||||||||
| SPEAR Base | 94M | 97k | 90.02 | 3.85 | 4.13 | 69.40 | 2.66 | 94.5 | 10.89 | 7.34 | 10.2 | 99.00 | ||
| SPEAR Large | 327M | 97k | 95.01 | 3.30 | 3.80 | 71.57 | 2.72 | 94.6 | 11.12 | 7.42 | 10.7 | 99.00 | ||
| SPEAR XLarge | 600M | 197k | 96.34 | 2.86 | 3.17 | 73.29 | 2.72 | 94.6 | 11.24 | 7.44 | 10.9 | 99.00 | ||
@@
As shown in Table 12 and Table 13, our speech-domain model SPEAR demonstrates very high performance on understanding, paralinguistics, and enhancement tasks on SUPERB. Our speech-domain model SPEAR Large achieves the same or better performance on 12 out of 15 tasks on SUPERB (except for ST, QbE, and VC) compared to WavLM Large, the previous state-of-the-art model on SUPERB, with the same pre-training corpora and similar model size. This again suggests that the performance of the SPEAR framework is not constrained by the teacher model, since SPEAR Large is pre-trained with fine-grained targets generated by WavLM Large. It has also been observed that performing dual-domain training leads to slight performance degradation on understanding and paralinguistic tasks compared to the speech-only models. However, improvement on KWS and enhancement tasks is also observed, suggesting a positive task synergy in these tasks. Finally, our largest dual-domain model, SPEAR XLarge, improves upon SPEAR Large, and further improves the best results of 12 SUPERB tasks, confirming that the SPEAR framework scales effectively with both model and data size.
It is worth noting that our results on the Voice Conversion (VC) task are not directly comparable to previous SSL models due to a change in the VC recipe within the SUPERB codebase, as pointed out by Shi et al. (2024). Within our own experiments, we observe that our Large and XLarge variants underperform the Base model on VC. This is potentially due to overfitting on the small training dataset, an issue also observed in MR-HUBERT (Shi et al., 2024). We leave the investigation of using our models for generation tasks as an important direction for future work, since a unified capability for both understanding and generation is highly desirable.
Appendix F HEAR Evaluation

### v2

@@ 5.3 HEAR Evaluation
Setup
To assess the general audio capabilities of our models, experiments are conducted on the HEAR benchmark (Turian et al., 2022), which evaluates audio representations across 19 diverse tasks. In contrast to SUPERB, HEAR focuses more on audio-related tasks requiring low-level acoustic details. The final-layer representations are used for evaluation unless otherwise specified. For clarity, the average scores for each of the three task categories are reported: Environment, Speech and Music, along with the overall average score in Table 5. More information regarding the tasks in HEAR and detailed results on individual tasks are provided in Appendix E.
Results
As shown in Table 5, all speech-domain models show limited overall performance on the audio-centric HEAR benchmark, highlighting the need to incorporate general audio data during pre-training. Nonetheless, our SPEAR models yield a higher overall score than the much larger WavLM Large, indicating that the fine-grained MVQ tokens capture low-level acoustic detail despite being extracted from a speech SSL model.
| Model | # Params | Pre-train Data | HEAR | |||
| Env | Speech | Music | Average | |||
@@ Results
| WavLM Base+ | 95M | 94k | 57.28 | 68.14 | 61.31 | 62.69 |
| WavLM Large | 317M | 94k | 72.86 | 72.69 | 65.77 | 69.65 |
@@
| Audio Models (AudioSet 5k hours) | ||||||
| BEATs | 90M | 5k | 73.23 | 62.40 | 77.52 | 71.05 |
| Dasheng-Base | 86M | 5k | - | - | - | 70.43 |
| Dasheng 0.6B | 600M | 5k | - | - | - | 71.75 |
| Dasheng 1.2B | 1.2B | 5k | - | - | - | 74.87 |
| Ours, SPEAR Base | 94M | 5k | 77.83 | 69.74 | 80.61 | 76.37 |
| Ours, SPEAR Large | 327M | 5k | 78.16 | 72.94 | 81.80 | 78.08 |
| Audio Models (others) | ||||||
| Dasheng-Base | 86M | 272k | 80.18 | 72.48 | 84.00 | 79.31 |
| Dasheng 0.6B | 600M | 272k | 82.95 | 74.82 | 84.73 | 81.03 |
| Dasheng 1.2B | 1.2B | 272k | 83.20 | 75.72 | 84.86 | 81.44 |
@@
| USAD Base | 94M | 126k | 80.67 | 73.72 | 79.31 | 77.75 |
| USAD Large | 330M | 126k | 81.97 | 74.48 | 81.7 | 79.36 |
| Ours, SPEAR Base | 94M | 97k | 80.60 | 73.95 | 79.35 | 77.83 |
@@
| Ours, SPEAR XLarge | 600M | 197k | 82.30 | 76.97 | 81.29 | 80.07 |
| Ours, SPEAR Base | 94M | 97k | 82.55 | 77.45 | 83.61 | 81.32 |
| Ours, SPEAR Large | 327M | 97k | 84.33 | 78.88 | 84.21 | 82.46 |
| Ours, SPEAR XLarge | 600M | 197k | 84.85 | 79.75 | 85.43 | 83.41 |
Among all audio SSL models trained on AudioSet, SPEAR achieves the highest average score on HEAR. Increasing pre-training data to 13k hours further improves the audio-domain SPEAR models. Notably, SPEAR Large achieves 83.58 on the environment category, surpassing the best performing audio model Dasheng 1.2B (83.2) pre-trained on 272k hours, with only a quarter of the model parameters and far less pre-training data. This highlights the potential of SPEAR for audio SSL at larger scales. However, the overall performance of SPEAR Large on HEAR still trails that of Dasheng 1.2B, a gap we attribute to the significantly smaller scale of general audio training data and model size.
@@
Our dual-domain SPEAR models consistently outperform their single-domain counterparts, verifying the benefits of unified pre-training over speech and audio data. The SPEAR Large achieves an average score of 79.26, surpassing both SPEAR Large and SPEAR Large, which suggests that joint speech–audio pre-training yields a more unified representation space. Compared to USAD Large, SPEAR Large pre-trained on a smaller corpora achieves an absolute improvement of 3.12 on the average score under the same evaluation setup of using intermediate representations. This again highlights the strength of SPEAR as a unified SSL framework for learning generic speech and audio representations. Compared to the much larger Dasheng 1.2B, SPEAR XLarge achieves stronger performance on speech-related tasks while trailing on environment and music tasks, which can be attributed to the imbalanced pre-training data composition (only 13k hours from the 197k hours used by SPEAR are general-audio data). However, it should be noted that SPEAR XLarge is a more versatile model than Dasheng 1.2B, outperforming Dasheng 1.2B on SUPERB (see Table 4) by a large margin.
5.4 Ablation Studies
Token Mixing
This ablation aims to investigate the effect of token mixing introduced in Section 3.2.3. By viewing WavLM-style denoising loss and our proposed token mixing mechanism effectively both as data augmentation methods, we compare three setups: without data augmentation, with WavLM-style loss (Chen et al., 2022), and with our proposed token mixing. The speech domain models are trained on LibriSpeech data using the Base scale with the aforementioned three configurations. Results are shown in Table 6. As can be seen, both augmentations outperform the baseline without augmentations. While achieving similar WERs on ASR, token mixing outperforms WavLM-style denoising on speaker diarisation (SD), speech separation (SS), and speech enhancement (SE) by a large margin, three tasks requiring separation from different sources in the overlapped/noisy speech. This validates the effectiveness of token mixing when processing complex sound scenes.
| Setup | LS-100 | SUPERB | |||||
| clean | other | PR | SD | SS | SID | SE | |
| None | 3.1 | 7.1 | 3.3 | 4.23 | 10.52 | 87.5 | 2.60 |
| WavLM-style | 3.1 | 7.0 | 3.4 | 3.78 | 10.76 | 88.4 | 2.63 |
| Token Mixing | 3.1 | 7.0 | 3.4 | 3.01 | 11.17 | 88.3 | 2.65 |
Dual-domain Training Strategy
As mentioned in Section 3.2.2, we adopt an asymmetrical pre-training loss for dual-domain pre-training. Here, we compare it with the other two strategies introduced in Section 3.2.2 using the dual-domain setup of Large size model on the SA-97k data (see Table 2). The models are evaluated after 100k training steps without token mixing augmentation. We compare the ASR and AT fine-tuning results, two SUPERB tasks, and the average score on HEAR, and the results are shown in Table 7. Among the three strategies, Disjoint achieves the worst performance, indicating that treating both domains independently is harmful to unified representation learning. Asymmetrical achieves the best overall performance, suggesting that learning speech MVQ tokens for audio data is a useful regularisation to bridge the domain mismatch. On the other hand, also computing pre-training loss against the audio MVQ tokens for speech data in the Joint strategy leads to slightly worse overall performance. One possible reason is that audio MVQ tokens contain limited semantic information relevant to speech, degrading the pre-training performance, especially when the speech data is dominant.
| Strategy | LS-100 | AS-20K | SUPERB | HEAR | ||
| clean | other | SID | PR | Avg | ||
| Joint | 2.9 | 5.9 | 36.9 | 90.6 | 3.34 | 78.7 |
| Disjoint | 3.0 | 5.8 | 37.0 | 87.4 | 3.40 | 78.3 |
| Asymmetrical | 2.9 | 5.8 | 36.9 | 90.7 | 3.12 | 79.0 |
Other Ablations and Comparisons
Beyond the studies above, we also investigated the effect of using different teacher models (Appendix F.1.1), encoder architectures (Zipformer vs Transformer, see Appendix F.3), discrete targets for pre-training (k-means vs MVQ, see Appendix F.4), the weighting factor in the masked-prediction pre-training loss (Appendix F.2), and various configurations of MVQ (Appendix F.5). A controlled comparison (same teachers, similar model size and training data) between USAD and SPEAR is conducted in Appendix G.
@@ 6 Conclusions
In this work, we propose SPEAR, a unified SSL framework for learning unified and generic representations across both speech and general audio domains. SPEAR leverages MVQ to generate fine-grained discrete speech and audio tokens from domain-specific teachers, and performs masked-token prediction for representation learning over speech and audio data. Using an asymmetrical dual-domain pre-training objective, SPEAR learns a balanced and unified representation space. Furthermore, a novel token mixing mechanism is designed to enhance the performance for complex sound scenes. The downstream fine-tuning experiments, along with the evaluation of frozen representations on two major benchmarks for assessing speech and general-audio representations (SUPERB and HEAR), demonstrate the effectiveness of SPEAR in learning unified and generic speech and audio representations. Our dual-domain model with 600M parameters excels in both domains, making it one of the most powerful and versatile open-source SSL models for unified speech and audio understanding.
Impact Statement
This work presents a unified representation model for both speech domain and general audio domain. It could serve as a unified frontend for various fields including speech understanding, voice modelling, sound event detection. This comprehensive audio perception capability could make the work an essential component for building multi-modal AI systems.
@@
Appendix A Multi-codebook Vector Quantisation
@@
A.1 Encode and Decode
A Multi-codebook Vector quantisation (MVQ) module consists of codebooks, each containing codebook vectors. Given a -dimensional representation , the MVQ quantiser encodes it to a sequence of integers (i.e, tokens) from a finite discrete value space .333For storage efficiency, we always use since the indices can be stored with uint8 format. However, it is theoretically possible to increase to a bigger number. The encoded integers are denoted as MVQ tokens, which can be used for reconstructing the original input through a operation:
@@ A.2 MVQ Training
Appendix B Model Specification and Training Settings
B.1 Model Specification
@@
B.2 Pre-training Resources and Configurations
The training hyperparameters and the required computing resources for training SPEAR are presented in Table 9. All models in SPEAR are trained using the NVIDIA A800 (80GB) GPUs. As shown in WavLM (Chen et al., 2022), applying data augmentations to the input audio can improve the pre-training performance. Therefore, we apply both in-batch utterance mixing and noise mixing during training. MUSAN (Snyder et al., 2015) is used as the noise dataset. The optimiser and scheduler settings follow (Yao et al., 2024), where the ScaledAdam optimiser and Eden scheduler are used. During pre-training, the Token Mixing is applied to 10% of the data, where the secondary signal is sampled from the same batch. During dual-domain training, we mix either speech or audio data into the speech training sample to simulate overlapped or noisy speech and perform token mixing using the MVQ tokens generated by the speech SSL teacher.
@@ B.2 Pre-training Resources and Configurations
| (see Equation 3.2.2) | - | - | - | - | 0.1 | 0.1 | 0.1 |
| Computing Resources | |||||||
| Num GPUs | 8 | 8 | 8 | 8 | 8 | 16 | 32 |
| GPU hours (approx.) | 460 | 900 | 290 | 560 | 660 | 2,000 | 3,800 |
B.3 Fine-tuning Configurations
The fine-tuning configurations for the downstream ASR tasks and AT tasks presented in Section 5.1 are shown in Table 10. We used Pruned RNN-T (Kuang et al., 2022), a memory-efficient variant of RNN-T for optimisation. An asynchronous learning rate policy is adopted during fine-tuning by setting a smaller learning rate for the pre-trained encoder parameters. In ASR experiments, a 3-fold speed perturbation is applied to the training data. In AS-2M, we follow prior work (Gong et al., 2021) to adopt a weighted sampler to cope with the imbalanced label distribution in the full AudioSet. Mixup (Zhang et al., 2018) with a probability of 0.5 is used in AT fine-tuning.
@@ B.3 Fine-tuning Configurations
Appendix C LibriSpeech Fine-tuning Experiments
@@
Appendix D SUPERB Evaluation
In this section, we provide more detail about the SUPERB benchmark (Yang et al., 2021; Tsai et al., 2022) as a supplement to Section 5.2 and present the complete results on the SUPERB benchmark. A summary of the SUPERB tasks is shown in Table 12. Complete results of SPEAR models along with other existing speech SSL models are shown in Table 13 and Table 14.
| Task Name | Metric(s) |
| Speaker Identification (SID) | Accuracy |
| Speaker Verification (SV) | Equal Error Rate (EER) |
| Speaker Diarization (SD) | Diarization error rate (DER) |
| Emotion Recognition (ER) | Accuracy |
| Phoneme Recognition (PR) | Phone Error Rate (PER) |
| Automatic Speech Recognition (ASR) | Word Error Rate (WER) |
| Out-of-domain ASR (ar/es/zh) | Word (character) Error Rate |
| Keyword Spotting (KS) | Accuracy |
| Query by Example Spoken Term Detection (QbE) | Maximum Term Weighted Value (MTWV) |
| Speech Translation (ST) | BLEU |
| Intent Classification (IC) | Accuracy |
| Slot Filling (SF) | F1, Character Error Rate (CER) |
| Speech Enhancement (SE) | Perceptual Evaluation of Speech Quality (PESQ) Short-Time Objective Intelligibility (STOI) |
| Speech Separation (SS) | Scale-invariant Signal-to-distortion Ratio improvement (SI-SDRi) |
@@ Appendix D SUPERB Evaluation
| Model | # Params | Pre-train Data | Understanding | ||||||||
| PR | ASR | OOD-ASR | KS | QbE | ST | IC | SF | ||||
| PER | WER | WER | Acc | MTWV | BLEU | Acc | F1 | CER | |||
| FBANK | 0 | - | 82.01 | 23.18 | 63.58 | 8.63 | 0.0058 | 2.32 | 9.10 | 69.64 | 52.94 |
| Existing Speech SSL models | |||||||||||
| WavLM Base+ (Chen et al., 2022) | 95M | 94k | 3.92 | 5.59 | 38.32 | 97.37 | 0.0988 | 24.25 | 99.00 | 90.58 | 21.20 |
| wav2vec 2.0 Large (Baevski et al., 2020) | 317M | 60k | 4.25 | 3.75 | 44.89 | 96.66 | 0.0480 | 12.48 | 95.28 | 87.11 | 27.31 |
| HuBERT Large (Hsu et al., 2021) | 317M | 60k | 3.53 | 3.62 | 44.08 | 95.29 | 0.0353 | 20.10 | 98.76 | 89.81 | 21.76 |
| WavLM Large (Chen et al., 2022) | 317M | 94k | 3.06 | 3.44 | 32.27 | 97.86 | 0.0886 | 26.57 | 99.31 | 92.21 | 18.36 |
| Ours, Speech SSL models | |||||||||||
| SPEAR Base | 94M | 84k | 3.34 | 3.43 | 34.20 | 97.50 | 0.0772 | 24.37 | 99.17 | 90.96 | 19.22 |
| SPEAR Large | 327M | 84k | 2.56 | 3.24 | 31.66 | 97.88 | 0.0768 | 26.20 | 99.44 | 92.25 | 17.86 |
| Ours, Speech & Audio SSL models | |||||||||||
| SPEAR Base | 94M | 97k | 3.84 | 3.79 | 35.37 | 97.58 | 0.0801 | 24.07 | 98.05 | 90.54 | 20.14 |
| SPEAR Large | 327M | 97k | 3.08 | 3.39 | 31.20 | 97.92 | 0.0712 | 25.64 | 99.40 | 92.07 | 18.04 |
| SPEAR XLarge | 600M | 197k | 2.92 | 3.15 | 30.51 | 98.12 | 0.0745 | 26.66 | 99.60 | 92.86 | 17.23 |
| Model | # Params | Pre-train Data | Paralinguistics | Enhancement | Generation | |||||||
| SID | SV | SD | ER | SE | SS | VC | ||||||
| Acc | EER | DER | Acc | PESQ | STOI | SI-SDRi | MCD | WER | SV | |||
| FBANK | 0 | - | 0 | 9.56 | 10.05 | 35.39 | 2.55 | 93.6 | 9.23 | 8.47 | 38.3 | 77.25 |
| Existing Speech SSL models | ||||||||||||
| WavLM Base+ (Chen et al., 2022) | 95M | 94k | 89.42 | 4.07 | 3.50 | 68.65 | 2.63 | 94.3 | 10.85 | 7.40 | 8.1 | 99.00 |
| wav2vec 2.0 Large (Baevski et al., 2020) | 317M | 60k | 86.14 | 5.65 | 5.62 | 65.64 | 2.52 | 94.0 | 10.02 | 7.63 | 15.8 | 97.25 |
| HuBERT Large (Hsu et al., 2021) | 317M | 60k | 90.33 | 5.98 | 5.75 | 67.62 | 2.64 | 94.2 | 10.45 | 7.22 | 9.0 | 99.25 |
| WavLM Large (Chen et al., 2022) | 317M | 94k | 95.49 | 3.77 | 3.24 | 70.62 | 2.70 | 94.5 | 11.19 | 7.30 | 9.9 | 99.00 |
| Ours, Speech SSL models | ||||||||||||
| SPEAR Base | 94M | 84k | 90.5 | 3.77 | 2.93 | 69.14 | 2.64 | 94.3 | 11.32 | 7.40 | 10.1 | 99.00 |
| SPEAR Large | 327M | 84k | 95.55 | 3.14 | 2.3 | 72.10 | 2.72 | 94.5 | 11.67 | 7.33 | 10.4 | 99.00 |
| Ours, Speech & Audio SSL models | ||||||||||||
| SPEAR Base | 94M | 97k | 90.26 | 3.85 | 3.0 | 69.24 | 2.67 | 94.5 | 11.39 | 7.34 | 10.2 | 99.00 |
| SPEAR Large | 327M | 97k | 95.19 | 3.30 | 2.31 | 71.66 | 2.73 | 94.6 | 11.60 | 7.42 | 10.7 | 99.00 |
| SPEAR XLarge | 600M | 197k | 96.34 | 2.86 | 1.99 | 73.29 | 2.75 | 94.6 | 11.73 | 7.44 | 10.9 | 99.00 |
@@
As shown in Table 13 and Table 14, our speech-domain model SPEAR demonstrates very high performance on understanding, paralinguistics, and enhancement tasks on SUPERB. Our speech-domain model SPEAR Large achieves better performance on 12 out of 15 tasks on SUPERB (except for ST, QbE, and VC) compared to WavLM Large, the previous state-of-the-art model on SUPERB, with the same pre-training corpora and similar model size. This again suggests that the performance of the SPEAR framework is not constrained by the teacher model, since SPEAR Large is pre-trained with fine-grained targets generated by WavLM Large. It has also been observed that performing dual-domain training leads to slight performance degradation on understanding and paralinguistic tasks compared to the speech-only models. However, improvement on KWS and enhancement tasks is also observed, suggesting a positive task synergy in these tasks. Finally, our largest dual-domain model, SPEAR XLarge, improves upon SPEAR Large, and further improves the best results of 12 SUPERB tasks, confirming that the SPEAR framework scales effectively with both model and data size.
It is worth noting that our results on the Voice Conversion (VC) task are not directly comparable to previous SSL models due to a change in the VC recipe within the SUPERB codebase, as pointed out by Shi et al. (2024a). Within our own experiments, we observe that our Large and XLarge variants underperform the Base model on VC. This is potentially due to overfitting on the small training dataset, an issue also observed in MR-HUBERT (Shi et al., 2024a). We leave the investigation of using our models for generation tasks as an important direction for future work, since a unified capability for both understanding and generation is highly desirable.
Appendix E HEAR Evaluation